# Semantics-Empowered Space-Air-Ground-Sea Integrated Network: New Paradigm, Frameworks, and Challenges

### v1

@@ Abstract
In the coming sixth generation (6G) communication era, to provide seamless and ubiquitous connections, the space-air-ground-sea integrated network (SAGSIN) is envisioned to address the challenges of communication coverage in areas with difficult conditions, such as the forest, desert, and sea. Considering the fundamental limitations of the SAGSIN including large-scale scenarios, highly dynamic channels, and limited device capabilities, traditional communications based on Shannon information theory cannot satisfy the communication demands. Moreover, bit-level reconstruction is usually redundant for many human-to-machine or machine-to-machine applications in the SAGSIN. Therefore, it is imperative to consider high-level communications towards semantics exchange, called semantic communications. In this survey, according to the interpretations of the term “semantics”, including “significance”, “meaning”, and “effectiveness-related information”, we review state-of-the-art works on semantic communications from three perspectives, which are 1) significance representation and protection, 2) meaning similarity measurement and meaning enhancement, and 3) ultimate effectiveness and effectiveness yielding. Sequentially, three types of semantic communication systems can be correspondingly introduced, namely the significance-oriented, meaning-oriented, and effectiveness/task-oriented semantic communication systems. Implementation of the above three types of systems in the SAGSIN necessitates a new perception-communication-computing-actuation-integrated paradigm (PCCAIP), where all the available perception, computing, and actuation techniques jointly facilitates significance-oriented sampling & transmission, semantic extraction & reconstruction, and task decision. Finally, we point out some future challenges on semantic communications in the SAGSIN. This survey provides a comprehensive review on the future semantic communications in the SAGSIN, and elaborates on the performance metrics and techniques related to semantic communications for references and further in-depth investigations.
@@ I-A Motivation
The sixth generation (6G) network has been envisioned to provide seamless and ubiquitous services in an extremely reliable and energy-efficient manner with ultra low latency[1, 2, 3]. However, the current widely used terrestrial network in fifth generation (5G) is facing severe challenges in facilitating such seamless and ubiquitous connections due to the following reasons. Firstly, with the proliferating number of connected devices and the surge of data traffic, the terrestrial network alone cannot satisfy such massive and broad connections. Secondly, almost 90 percent of the Earth area is covered by the sea, forest, and desert, where the terrestrial network cannot provide continuous connection for its harsh environment. To tackle the above inherent challenges in the current 5G network, terrestrial networks should be extended by integrating the nodes deployed in the space, air, and (or) sea to form a space-air-ground(-sea) integrated network (SAGIN or SAGSIN), which has garnered a surge of interest by researchers[4, 5, 6, 7, 8, 9, 10, 11]. The SAGSIN is comprised of space-based network, air-based network, and sea-based network along with ground-based (terrestrial) network, as shown in Fig. 1. Specifically, space-based network consists of geostationary Earth orbit (GEO) satellites and low Earth orbit (LEO) satellites, which usually form constellations respectively; air-based network is composed of airplanes, unmanned aerial vehicles (UAVs), and high-altitude platforms (HAPs), which usually appear as platoons; sea-based network includes nodes such as maritime base stations, ships, and unmanned ships; ground-based network has been extensively developed with the advent of cell mobile communications, autonomous driving, and Internet of Things (IoT). By connecting the nodes within a certain part via short or long wireless links to construct the abovementioned four types of networks, and integrating the four parts via inter-network links, the SAGSIN can perfectly reap the promised benefits of 6G communications owing to its massive connections among long-distance nodes. However, the realization of 6G communications with extremely high reliability, very low latency, and high energy efficiency based on the SAGSIN will be affected by several fundamental limitations, which are large-scale scenarios, highly dynamic channels, and limited device capabilities. Firstly, the communication distance between the transceiver nodes in the SAGSIN scenarios is considerably large in stark comparison with terrestrial communications. For instance, the distance from LEO to ground base station is 500-2,000 kilometers, which causes inherent and non-negligible propagation latency. Secondly, the communication channels over transceivers in the SAGSIN vary rapidly compared with terrestrial channels. Taking UAV-LEO-ground multi-hop communications as an example, the LEO orbits around the Earth with significantly high velocity, leading to inherent Doppler shift in UAV-LEO and LEO-ground channels; besides, cosmic environment is so harsh that deep fading will appear due to cosmic ray and atmospheric attenuation. Thirdly, the available communication and computing resources are strictly limited for node devices in the SAGSIN. Specifically, due to hardware and battery limitation on the satellites, UAVs, and so on, transmission rate and computing speed of space/air-ground communications are remarkably lower in comparison to terrestrial communications. The abovementioned inherent and fundamental characteristics of the SAGSIN impose severe challenge on current traditional communication technologies in facilitating the envisioned 6G communications. On the other hand, traditional communications based on classic Shannon information theory are reaching their limits, which may not support communications in the SAGSIN with such severe limitations. Concretely speaking, since Shannon completed his masterpiece which lays the foundation of information theory in 1948[12], researchers have contributed much effort to find state-of-the-art coding and modulation schemes that can reach the channel capacity. From near Shannon limit codes such as Turbo codes[13] and low density parity check (LDPC) codes[14], to capacity-achievable channel codes such as Polar codes[15] and Spinal codes[16], the Shannon limit has been reached approximately by these advanced channel coding techniques. From frequency division multiple access (FDMA) applied in the first generation (1G) communications, to non-orthogonal multiple access (NOMA) emerging in 5G communications, the revolution of modulation methods has also enhanced the channel capacity and been reaching the Shannon limits. Therefore, a new revolution on current communication technologies is imperative to satisfy the proliferating need on data exchange in the SAGSIN.
Moreover, in many human-to-machine and machine-to-machine application scenarios in the SAGSIN, the ultimate goal of communications is to complete specific tasks at the machine terminal. Thus, perfect or approximate reconstruction of information on bit level is usually unnecessary. For instance, in the ship detection tasks, the observer can detect the ships based on only the pixels of image related to ships, where reconstruction of the whole image is evidently redundant. Another example is emergency monitoring and rescue tasks, where monitors are expected to only perceive and transmit urgent status information, while missed detection or error transmission of other statuses that are not very urgent are tolerable. Under harsh conditions of the SAGSIN, pursuing perfect or approximate restoration of raw data not only will consume a huge amount of resources but is also not feasible due to limited device capabilities. Therefore, the above reasons necessitate an idea change from traditional communication technologies to high-level communications towards semantics exchange, named semantic communications[17], for semantic communications aim at recovering only semantics at the receiver instead of reconstructing every code symbol accurately. In [17], Weaver identifies the three levels of communications, which are technical communications, semantic communications, and effectiveness communications. The abovementioned state-of-the-art coding and modulation schemes all make remarkable contributions to technical communications, while deliberately neglecting the semantic and effectiveness levels of communications, since Shannon believes that the technology of symbol transmission can be independent from the semantics that the symbols hold. However, the idea change from technical communications to semantic communications inspires us to reconsider the assumption of semantic independence and take semantic aspects of communications into consideration. This serves as the principal line of this survey.111It is worth noting that in this survey, we consider both semantic level and effectiveness level communications as semantic communications. This survey is dedicated to answering the following three questions regarding to semantic communication systems deployed in the SAGSIN. Question 1. How should we interpret ”semantics” for semantic communications? Question 2. How should we design semantic communication systems based on the interpretations of “semantics”? Question 3. How should we utilize the available while limited resources in the SAGSIN to implement semantic communications? Overall, the term “semantics” has three interpretations in state-of-the-art research on semantic communications. Firstly, from the perspective of etymology, the word “semantics” originates from the Greek “sēmanticós”, which means “significance”. Secondly, “semantics” has its common explanation, which is “meaning”. Thirdly, since “semantic communications” includes both semantic and effectiveness level, this term can also be interpreted as “effectiveness-related information”. Based on above, we introduce three types of semantic communication systems, which are respectively significance-oriented, meaning-oriented, and effectiveness-oriented semantic communication systems. (1) Significance-oriented semantic communication system. In fact, source data are not equally significant, and thus sampling and transmitting unimportant data will cause waste in resource consumption. Taking the emergency monitoring task as an example, the status message generated later is more significant than the one generated earlier, since the former contains fresher information for the receiver; moreover, the status message which is different from the previous one is more significant, since the status variation usually implies policy or decision change for the task actuator; besides, an abnormal status message is more significant than a normal one, since the environment becomes more urgent due to potential severe consequences. Therefore, from the significance perspective, source data should be sampled, coded, and modulated in a significance-oriented manner, such that only data with more semantics (i.e., more significant data) are reserved and transmitted while those with less semantics (i.e., less significant data) get discarded. How we can measure the “significance” will be elaborated in Section III. (2) Meaning-oriented semantic communication system. After significance-oriented sampling, although the amount of data has been preliminarily reduced, the remaining sampled data still include considerably large amount of redundancy due to semantics-unrelated information. Therefore, semantic extraction from the sampled data should be further adopted at the transmitter in order to eliminate the meaning-unrelated redundancy, such that the receiver can reconstruct data with similar meaning to original data. In the era of Shannon, semantic extraction and reconstruction were tough problems, and researchers had to temporarily neglect the semantic aspects of source data and dedicate themselves to technical communication revolution. Nowadays, the thriving of artificial intelligence (AI), especially the unprecedented proliferation of deep learning (DL) applications, will enable deep feature (semantics) extraction by massive computing resources, and utilizing DL to realize semantic communication systems is becoming a major approach. How DL-based semantic communication systems are designed will be discussed in Section IV. (3) Effectiveness-oriented semantic communication system. After extraction of meaning-related information from sampled data, most of the redundancy is eliminated with only the meaning of data to be transmitted. However, not all of the extracted meaning is useful for the receiver, since there is still redundant information for the ultimate actuation. Moreover, symbol- or semantics-level reconstruction is also unnecessary from effectiveness-oriented perspective, since task actuation process may be intelligently conducted based directly on received messages. Therefore, a task-oriented design for semantic communications can be adopted in order that the sampled data yield ultimate effectiveness more efficiently at the terminal. Specifically, a more intelligent effectiveness-related semantic extractor is introduced to replace common semantic extractor for capturing effectiveness-related information. Moreover, taking full advantage of DL-based techniques, intelligent actuator is adopted which receives the effectiveness-related information and directly outputs the actuation results of the task without reconstruction of original data. By such task-oriented design, the reconstruction process at the receiver gets omitted and thus computing resource consumption will decline. How such task-oriented communication system is constructed can be found in Section V. Furthermore, in order to implement semantic communication systems in the SAGSIN from the above three perspectives, all the available perception, computing, and actuation techniques should be jointly adopted. Specifically, significance-oriented semantic communication system adopts a perception-communication-integrated design by filtering insignificant information and protecting significant one; meaning-oriented semantic communication system utilizes perception, communication, and computing techniques through DL-based meaning extraction and reconstruction; effectiveness-oriented semantic communication system integrates all the perception, communication, computing, and actuation techniques via effectiveness-related semantic extractor and intelligent actuator without reconstruction. Therefore, a new paradigm called Perception-Communication-Computing-Actuation Integrated Paradigm (PCCAIP) can be proposed that guides the design of the above three types of semantic communication systems in the SAGSIN.
I-B Related Works
@@
Given that there is a recent surge of research interest on semantic communications, quite a few review and tutorial works on this topic have also emerged in the literature. In the aspects of short briefs, [18] envisions a significance-oriented semantic communication system with sparse and effectiveness-aware sampling, where semantics of data is measured by timeliness metrics and end-to-end mean square error (MSE). Another brief review [19] constitutes a tutorial on the framework of semantic communications by reconsidering semantic information theory. From semantic network perspective, [20] proposes a semantic-aware network architecture by utilizing federated edge learning. Moreover, [21] presents a comprehensive review on DL-based semantic communications. Nevertheless, all these brief reviews on semantic communications focus on only a certain interpretation of “semantics” listed in the previous subsection.
@@
Researchers have also contributed comprehensive tutorial works on semantic communication topics. The first of such tutorial works is [22], where three communication modalities including human-human, human-machine, and machine-machine are addressed by semantic communications using DL-based techniques. In [23], semantic communication theory is reviewed, and DL-based semantic communications for different types of sources are respectively discussed. In [24], a new perspective called reasoning-driven semantic communication system is proposed, and a new semantic representation method called semantic language is studied for next generation communications. Moreover, in the recent literature of [25], the authors review semantic communications from perspectives of semantic measures, semantic compression based on semantic rate-distortion theory, semantic transmission via joint source-channel coding (JSCC), and timeliness of semantic communications, respectively. In the recent tutorial work [26], research on semantic communications is reviewed from three perspectives, which are semantic-oriented communications, goal-oriented communications, and semantic-awareness communications, and correspondingly three directions on semantic communications are pointed out, which are semantic extraction, semantic transmission, and semantic metrics. In addition to the above tutorial works, some other tutorials on edge learning[27] and new challenges of 6G[28] also consider semantic communications as competitive solutions. However, these studies does not provide an integrated paradigm of semantic communication system design that satisfies the needs in the SAGSIN.
I-C Contributions

### v3

@@ Abstract
In the coming sixth generation (6G) communication era, to provide seamless and ubiquitous connections, the space-air-ground-sea integrated network (SAGSIN) is envisioned to address the challenges of communication coverage in areas with difficult conditions, such as the forest, desert, and sea. Considering the fundamental limitations of the SAGSIN including large-scale scenarios, highly dynamic channels, and limited device capabilities, traditional communications based on Shannon information theory cannot satisfy the communication demands. Moreover, bit-level reconstruction is usually redundant for many human-to-machine or machine-to-machine applications in the SAGSIN. Therefore, it is imperative to consider high-level communications towards semantics exchange, called semantic communications. In this survey, according to the interpretations of the term “semantics”, including “significance”, “meaning”, and “effectiveness-related information”, we review state-of-the-art works on semantic communications from three perspectives, which are 1) significance representation and protection, 2) meaning similarity measurement and meaning enhancement, and 3) ultimate effectiveness and effectiveness yielding. Sequentially, three types of semantic communication systems can be correspondingly introduced, namely the significance-oriented, meaning-oriented, and effectiveness/task-oriented semantic communication systems. Implementation of the above three types of systems in the SAGSIN necessitates a new perception-communication-computing-actuation-integrated paradigm (PCCAIP), where all the available perception, computing, and actuation techniques jointly facilitate significance-oriented sampling & transmission, semantic extraction & reconstruction, and task decision. Finally, we point out some future challenges on semantic communications in the SAGSIN. This survey provides a comprehensive review on the future semantic communications in the SAGSIN, and elaborates on the performance metrics and techniques related to semantic communications for references and further in-depth investigations.
@@ I-A Motivation
The sixth generation (6G) networks have been envisioned to provide seamless and ubiquitous services in an extremely reliable and energy-efficient manner with ultra low latency[1, 2, 3]. Especially, heterogeneous communication nodes distributed in city, rural area, no man’s land, as well as sea, air, and space, are envisioned to be seamlessly and continuously connected via 6G networks. However, the widely used terrestrial networks in fifth generation (5G) communications are facing severe challenges in facilitating such seamless and ubiquitous connections due to the following reasons. Firstly, with the proliferating number of connected devices, the surge of data traffic, and the rapidly growing demand of communication coverage, the terrestrial network alone cannot satisfy such massive and broad connections due to limited communication distance of terrestrial 5G networks. Secondly, almost 90 percent of the Earth area is covered by the sea, forest, and desert where no people live, and the 5G terrestrial networks are not accessible in these areas for the harsh environment and expensive infrastructure costs. To tackle the above inherent challenges in the current 5G network, terrestrial networks should be extended by integrating the nodes deployed in the space, air, and (or) sea to form a space-air-ground(-sea) integrated network (SAGIN or SAGSIN), which has garnered a surge of interest by researchers[4, 5, 6, 7, 8].
The SAGSIN is comprised of space-based network, air-based network, and sea-based network along with ground-based (terrestrial) network, as shown in Fig. 1. Specifically, space-based network consists of geostationary Earth orbit (GEO) satellites and low Earth orbit (LEO) satellites, which usually form constellations respectively; air-based network is composed of airplanes, unmanned aerial vehicles (UAVs), and high-altitude platforms (HAPs), which usually appear as platoons; sea-based network includes nodes such as maritime base stations, ships, and unmanned ships; ground-based network has been extensively developed with the advent of cell mobile communications, autonomous driving, and Internet of Things (IoT). By connecting the nodes within a certain part via short or long wireless links to construct the abovementioned four types of networks, and integrating the four parts via inter-network links, the SAGSIN can perfectly reap the promised benefits of 6G communications owing to its massive connections among long-distance nodes.
However, the realization of 6G communications with extremely high reliability, very low latency, and high energy efficiency based on the SAGSIN will be affected by several fundamental limitations, which are large-scale scenarios, highly dynamic channels, and limited device capabilities. Firstly, the communication distance between the transceiver nodes in the SAGSIN scenarios is considerably large in stark comparison with terrestrial communications. For instance, the distance from LEO satellites to ground nodes (e.g. base stations or user equipment) is 500-2,000 kilometers, which causes inherent and non-negligible propagation latency. Secondly, the communication channels over transceivers in the SAGSIN vary rapidly compared with terrestrial channels. Taking UAV-LEO-ground multi-hop communications as an example, the LEO orbits around the Earth with significantly high velocity, leading to inherent Doppler shift in UAV-LEO and LEO-ground channels; besides, cosmic environment is so harsh that deep fading will appear due to cosmic ray and atmospheric attenuation. Thirdly, the available communication and computing resources are strictly limited for node devices in the SAGSIN. Specifically, due to hardware and battery limitation on the satellites, UAVs, and so on, transmission rate and computing speed of space/air-ground communications are remarkably lower in comparison to terrestrial communications.
The abovementioned inherent and fundamental characteristics of the SAGSIN impose severe challenges on current traditional communication technologies in facilitating the envisioned 6G communications. On the other hand, traditional communications based on classic Shannon information theory are reaching their limits, which may not support communications in the SAGSIN with such severe limitations. Concretely speaking, since Shannon completed his masterpiece which lays the foundation of information theory in 1948[9], researchers have contributed much effort to find state-of-the-art coding and modulation schemes that can reach the channel capacity. From near Shannon limit codes such as Turbo codes[10] and low density parity check (LDPC) codes[11], to capacity-achievable channel codes such as Polar codes[12] and Spinal codes[13], the Shannon limit has been reached approximately by these advanced channel coding techniques. From frequency division multiple access applied in the first generation communications, to non-orthogonal multiple access (NOMA) emerging in 5G communications, the revolution of modulation methods has also enhanced the channel capacity and been reaching the Shannon limits. Therefore, a new revolution on current communication technologies is imperative to satisfy the proliferating need on data exchange in the SAGSIN.
Moreover, in many human-to-machine and machine-to-machine application scenarios in the SAGSIN, the ultimate goal of communications is to complete specific tasks at the machine terminal. Thus, perfect or approximate reconstruction of information on bit level is usually unnecessary. For instance, in the ship detection tasks, the observer can detect the ships from an image based on only the pixels related to ships, where reconstruction of the whole image is evidently redundant. Another example is emergency monitoring and rescue tasks, where monitors are expected to only perceive and transmit urgent status information, while missed detection or error transmission of other statuses that are not very urgent are tolerable. Under harsh conditions of the SAGSIN, pursuing perfect or approximate restoration of raw data not only will consume a huge amount of resources but is also not feasible due to limited device capabilities.
Therefore, the above reasons necessitate an idea change from traditional communication technologies to high-level communications towards semantics exchange, named semantic communications[14]. The reason is that semantic communications aim at recovering only semantics at the receiver instead of reconstructing every code symbol accurately. In [14], Weaver identifies the three levels of communications, which are technical communications, semantic communications, and effectiveness communications. The abovementioned state-of-the-art coding and modulation schemes all make remarkable contributions to technical communications, while deliberately neglecting the semantic and effectiveness levels of communications. This is because Shannon believes that the technology of symbol transmission can be independent from the semantics that the symbols hold. However, the idea change from technical communications to semantic communications inspires us to reconsider the assumption of semantic independence and take semantic aspects of communications into consideration. This serves as the principal line of this survey. That is, by means of semantic and effectiveness communications, the semantics of the source data can be extracted at the transmitter side and utilized for task actuation at the receiver side. 111It is worth noting that in this survey, we consider both semantic and effectiveness communications as semantic communications, since they both conduct data pre-processing (i.e., extraction of semantics) before data transmission. According to [14], the principal difference of semantic and effectiveness communication is whether the effectiveness of terminal task actuation is considered in system design, and this is why we review task-oriented communications (a kind of effectiveness communications) in Section V.
This survey is dedicated to answering the following three questions regarding to semantic communication systems deployed in the SAGSIN.
Question 1. How should we interpret “semantics” for semantic communications?
Question 2. How should we design semantic communication systems based on the interpretations of “semantics”?
Question 3. How should we utilize the available while limited resources in the SAGSIN to implement semantic communications?
Overall, the term “semantics” has three interpretations in state-of-the-art research on semantic communications. Firstly, from the perspective of etymology, the word “semantics” originates from the Greek “sēmanticós”, which means “significance”. Secondly, “semantics” has its common explanation, which is “meaning”. Thirdly, since “semantic communications” includes both semantic and effectiveness level, this term can also be interpreted as “effectiveness-related information”.
Based on above, we introduce three types of semantic communication systems, which are respectively significance-oriented, meaning-oriented, and effectiveness-oriented semantic communication systems.
(1) Significance-oriented semantic communication system. In fact, source data are not equally significant, and thus sampling and transmitting unimportant data will cause waste in resource consumption. Taking the emergency monitoring task as an example, the status message generated later is more significant than the one generated earlier, since the former contains fresher information for the receiver; moreover, the status message which is different from the previous one is more significant, since the status variation usually implies policy or decision change for the task actuator; besides, an abnormal status message is more significant than a normal one, since urgent abnormal statuses may cause potential severe consequences. Therefore, from the significance perspective, source data should be sampled, coded, and modulated in a significance-oriented manner, such that only data with more semantics (i.e., more significant data) are reserved and transmitted while those with less semantics (i.e., less significant data) get discarded. How we can measure the “significance” will be elaborated in Section III.
(2) Meaning-oriented semantic communication system. After significance-oriented sampling, although the amount of data has been preliminarily reduced, the remaining data still include considerably large amount of redundancy due to meaning-unrelated information. Therefore, semantic extraction from the data should be further adopted at the transmitter in order to eliminate the meaning-unrelated redundancy, such that the receiver can reconstruct data with similar meaning to original data. In the era of Shannon, semantic extraction and reconstruction were tough problems, and researchers had to temporarily neglect the semantic aspects of source data and dedicate themselves to technical communication revolution. The principal reason is that the semantics of different modals of source data are distinct and complex, which cannot be mathematically expressed in a unified form. Nowadays, the thriving of artificial intelligence (AI), especially the unprecedented proliferation of deep learning (DL) applications, will enable deep feature (semantics) extraction from various data such as texts, images, audios, and videos by massive computing resources, and utilizing DL to realize semantic communication systems is becoming a major approach. How DL-based semantic communication systems are designed for various modals of data will be discussed in Section IV.
(3) Effectiveness-oriented semantic communication system. After extraction of meaning-related information from the data, most of the redundancy is eliminated with only the meaning of data to be transmitted. However, not all of the extracted meaning is useful for the receiver, since there is still redundant information for the ultimate actuation. Moreover, symbol- or semantics-level reconstruction is also unnecessary from effectiveness-oriented perspective, since task actuation process may be intelligently conducted based directly on received messages. Therefore, a task-oriented design for semantic communications can be adopted in order that the source data yield ultimate effectiveness more efficiently at the terminal. Specifically, a more intelligent effectiveness-related semantic extractor is introduced to replace common semantic extractor for capturing effectiveness-related information. Moreover, taking full advantage of DL-based techniques, intelligent actuator is adopted which receives the effectiveness-related information and directly outputs the actuation results of the task without reconstruction of original data. By such task-oriented design, the reconstruction process at the receiver gets omitted and thus computing resource consumption will decline. How such task-oriented communication system is constructed can be found in Section V.
Furthermore, in order to implement semantic communication systems in the SAGSIN from the above three perspectives, all the available perception, computing, and actuation techniques should be jointly adopted. Specifically, significance-oriented semantic communication system adopts a perception-communication-integrated design by filtering insignificant information and protecting significant one; meaning-oriented semantic communication system utilizes perception, communication, and computing techniques through DL-based meaning extraction and reconstruction; effectiveness-oriented semantic communication system integrates all the perception, communication, computing, and actuation techniques via effectiveness-related semantic extractor and intelligent actuator without reconstruction. Therefore, a new paradigm called Perception-Communication-Computing-Actuation Integrated Paradigm (PCCAIP) can be proposed that guides the design of the above three types of semantic communication systems in the SAGSIN.
I-B Related Surveys & Tutorials on SAGSIN
Since implementing SAG(S)IN in future 6G networks has become as a consensus from both industry and academia, there have been a plethora of survey papers on the topic of SAG(S)IN. Specifically, many magazine articles have contributed to the tutorial works on various applications facilitated by SAG(S)IN. [15, 16] focuses on random access technologies in space communications as well as SAGIN-based communications. In [17], challenges and opportunities of non-terrestrial network (including space, air, and sea networks in SAGSIN) in 6G era are elaborated through a case study on the millimeter wave communication among air/space nodes and ground nodes. The authors in [18] also give a brief tutorial on new three-dimension communication architecture through SAGSIN. In terms of newly emerging technologies such as software-defined radio[19, 20], optical communications[21], and radar communications[22], SAG(S)IN is also recommended as promising application scenarios. Recently, with the surge of AI applications, optimizing the performance of next generation communications facilitated by SAGIN has raised much attention[23, 24, 25, 26, 27]. Moreover, some tutorial works on simulation platforms of SAGIN scenario can guide the communication system design of future 6G network[28].
There are also some comprehensive survey papers that elaborate on the characteristics, design methods, and challenges of implementing SAG(S)IN. The first of such survey papers is [4], where the issues of system integration, protocol design, and performance analysis & optimization are discussed in the integrated networks in order to overcome the heterogeneity (i.e., limited device capabilities), self-organization, and time-variability (i.e., highly dynamic channels) characteristics of SAGIN. Inspired by this pioneering work, [5] gives a tutorial on how SAGIN facilitates next-generation communications in high-speed railway scenarios. Moreover, in order to address the security issues of heterogeneous SAGIN, blockchain-empowered space-air-ground IoT is proposed and the state-of-the art solutions for implementing blockchain-empowered SAGIN are reviewed in [6]. Security issues in SAGSIN are also the focus of [7], where a comprehensive tutorial is presented on the topic of secure threats as well as attack technologies and corresponding defense methods. Also, the recent work [8] envisions the integration of terrestrial network and non-terrestrial network in the future 6G era.
However, the above tutorial works all pay close attention to only traditional communication technologies. As has been emphasized before, the inherent characteristics of SAGSIN may hinder the implementation of bit-level communications because of Shannon limits. Moreover, perfect bit reconstruction will be not that necessary in human-to-machine (or machine-to-machine) communications in future 6G networks. Hence, an idea change from traditional bit-level communication to semantic-level communication is imperative, especially in the future SAGSIN.
I-C Related Surveys & Tutorials on Semantic Communications
@@
Given that there is a recent surge of research interest on semantic communications, quite a few review and tutorial works on this topic have also emerged in the literature. In [29], the authors envision a significance-oriented semantic communication system with sparse and effectiveness-aware sampling, where semantics of data is measured by timeliness metrics and end-to-end mean square error (MSE). Another magazine article [30] gives a tutorial on the framework of semantic communications by reconsidering semantic information theory. From the perspective of semantic network, [31] proposes a semantic-aware network architecture by utilizing federated edge learning. Moreover, [32] presents a comprehensive review on DL-based semantic communications. Nevertheless, all these articles on semantic communications focus on only a certain interpretation of “semantics” listed in the previous subsection.
Survey Papers | Contributions | |||
Related to SAGSIN | Magazine Articles | [15, 16] | Random access technologies in SAGSIN | |
[17] | Non-terrestrial network for 6G | |||
[18] | Three-dimension communication architecture of SAGSIN | |||
[19, 20] | Software-defined radio facilitated by SAGSIN | |||
[21, 22] | Optical communications and radar communications in SAGSIN | |||
[23, 24, 25, 26, 27] | AI-enabled SAGSIN | |||
[28] | Simulation platform establishment for SAGSIN | |||
Comprehensive Surveys | [4] |
[5] | Tutorial on high-speed railway facilitated by SAGSIN | |||
[6] | Blockchain-empowered SAGSIN aiming at security issues | |||
[7] | SAGSIN security issues including attack & defense methods | |||
[8] | Integrating terrestrial network and non-terrestrial network in 6G era | |||
Related to Semantic Communications | Magazine Articles | [29] | Significance-oriented communications via effectiveness-aware sampling | |
[30] | Reconsidering semantic information theory | |||
[31] | Federated edge learning-based semantic-aware network | |||
[32] | Comprehensive review on DL-based semantic communications | |||
Comprehensive Surveys | [33] |
[34] |
[35] |
[36] |
[37] | Semantic-oriented, goal-oriented, and semantic-aware communications | |||
[38, 39] | Semantic communications as solutions in future networks | |||
Our Survey | Contributions for SAGSIN | Contributions for Semantic communications | ||
Proposes PCCAIP as a | Reviewing SPTs for significance-oriented communications | |||
unified framework facilitating | Reviewing METs for meaning-oriented communications | |||
task-critical applications | Reviewing EYTs for effectiveness-oriented communications |
@@
Some comprehensive tutorial works on semantic communication topics can also be found in the literature. The first of such tutorial works is [33], where three communication modalities including human-human, human-machine, and machine-machine are addressed by semantic communications using DL-based techniques. In [34], semantic communication theory is reviewed, and DL-based semantic communications for different types of sources are respectively discussed. In [35], a new perspective called reasoning-driven semantic communication system is proposed, and a new semantic representation method called semantic language is studied for next generation communications. Moreover, in the recent literature of [36], the authors review semantic communications from perspectives of semantic measures, semantic compression based on semantic rate-distortion theory, semantic transmission via joint source-channel coding (JSCC), and timeliness of semantic communications, respectively. In the recent tutorial work [37], research on semantic communications is reviewed from three perspectives, which are semantic-oriented communications, goal-oriented communications, and semantic-awareness communications. Correspondingly, the three directions on semantic communications are pointed out, which are semantic extraction, semantic transmission, and semantic metrics. In addition to the above tutorial works, some other tutorials on edge learning[38] and new challenges of 6G[39] also consider semantic communications as competitive solutions.
However, a principal implementation of future 6G network is SAGSIN, while these studies do not provide an integrated paradigm of semantic communication system design that satisfies the needs in the SAGSIN. For instance, communication over highly dynamic channels is a core characteristic of SAGSIN, and thus robust semantic communications that resist channel fading as well as semantic noise must be taken into consideration when implementing semantic communications in SAGSIN. Also, given the limited capabilities of nodes in SAGSIN, it is crucial to prioritize semantic rate allocation and reduce computing complexity. Moreover, as has been pointed out in [6, 7], security issues should be addressed in SAGSIN-based semantic communications. Goal-oriented communications for specific task-critical applications in SAGSIN are also in its nascent age. As a result, a comprehensive review of semantic communications in the future SAGSIN is of vital importance.
I-D Contributions